OpenStudy (anonymous):

how will you find the slope of the line in 3x-2y+2=0 ?

OpenStudy (earthcitizen):

y=(3x+2)/2 m=3/2

OpenStudy (anonymous):

\(3x-2y+2=0\) is in standard form, you have to convert it in the slope intercept form. \(y=\frac32x+1 \) hence the slope is \(y=\frac32\)

OpenStudy (anonymous):

thanks :)

OpenStudy (anonymous):

glad to help :)

OpenStudy (anonymous):

another thing, can you show your solution for that ?